# Modeling Unknown Stochastic Dynamical System Subject to External Excitation

### v1

@@ Abstract
We present a numerical method for learning unknown nonautonomous stochastic dynamical system, i.e., stochastic system subject to time dependent excitation or control signals. Our basic assumption is that the governing equations for the stochastic system are unavailable. However, short bursts of input/output (I/O) data consisting of certain known excitation signals and their corresponding system responses are available. When a sufficient amount of such I/O data are available, our method is capable of learning the unknown dynamics and producing an accurate predictive model for the stochastic responses of the system subject to arbitrary excitation signals not in the training data. Our method has two key components: (1) a local approximation of the training I/O data to transfer the learning into a parameterized form; and (2) a generative model to approximate the underlying unknown stochastic flow map in distribution. After presenting the method in detail, we present a comprehensive set of numerical examples to demonstrate the performance of the proposed method, especially for long-term system predictions.
@@ 1 Introduction
There has been a growing interest in recovering/discovering unknown dynamical systems from observational data. Most of the existing studies focus on deterministic systems, with methods such as physics-informed neural networks (PINNs) ([39, 40]), SINDy ([5]), Fourier neural operator (FNO) ([24]), computational graph completion ([30]), sparsity promoting methods ([41, 42, 19]), flow map learning (FML) ([38, 11]), to name a few.
@@
Learning unknown stochastic systems is notably more challenging, as the stochastic noises in the systems usually can not be directly observed. The existing work utilizes Gaussian process ([48, 1, 12, 29]), polynomial approximations ([44, 22]), deep neural networks (DNNs) [8, 47, 9, 49, 14, 50], etc. More recently, a stochastic extension of the deterministic flow map learning (FML) approach ([38, 11]) was proposed. It employs generative models such as GANs (generative adversarial networks) ([10]) or autoencoders ([46]) to model the underlying stochasticity. However, most, if not all, of these methods are developed for autonomous systems, where time-invariance (in distribution) holds true and is critical to the method development.
@@
The focus and contribution of this paper is on the learning and modeling of unknown non-autonomous stochastic systems. More specifically, we consider SDEs with unknown governing equations and subject to time dependent external excitation or control signals. Our goal is to develop a method that can capture the stochastic dynamics of the unknown systems by using short-term data consisting of input/output (I/O) relations between excitation signals and their corresponding system responses. We remark that there exist some studies on modeling deterministic non-autonomous systems, using methodology such as Dynamic Mode Decomposition (DMD) ([25, 34]), SINDy ([6]), Koopman operator ([35]), FML ([36]), etc. These methods are not applicable for stochastic non-autonomous systems.
@@
The proposed method in this paper has two key components. First, the method utilize the observational I/O data to construct an accurate representation of unknown stochastic dynamics of the system. This is accomplished by a generative model that learns the stochastic mapping of the system between two consecutive discrete time steps. The learning of this stochastic flow map is similar to the work of [10, 46], which extended the deterministic FML to stochastic systems. While [10, 46] utilized GANs and autoencoder as the generative model, in this paper we employ conditional normalizing flow (cf. [31]). Normalizing flow has been widely adopted as a probabilistic model for generating data with desired distributions. Its applications include image and video generation [21], statistical inference and sampling [27, 43], reinforcement learning [18], as well as scientific computing [26, 23, 17, 13]. The second key component of the proposed method is local parameterization of the excitation signal in the training I/O data. The method was first introduced in [36] for deterministic nonautonomous sytem. We adopt the similar idea and extend it to stochastic system. The approach seeks to parameterize the excitation sigals in the training data via a localized polynomial over one time step. This then transforms the learning problem into a parametric learning between the coefficients of the local polynomials and the system responses. This is a critical component, as it allows the learned system to conduct long-term system predictions under arbitrary excitation signals that are never seen in the training data. Although the proposed method requires a large number of short bursts data, the overall demand for data may not be as large. This is because the burst length of the training I/O data is as short as two time steps. Once trained, the learned model is able to simulate the unknown stochastic systems for very long-term and subject to arbitrary exitation/control signals. We demonstrate this important property in several of our numerical examples. The learning is performed using training I/O data observed over only $O(1)$ nondimensional time units. However, the system predictions by the learned system can be accurate for time units as large as $O(10^{4})$ and beyond, and under excitation signals not in the training data.
@@ 2 Setup
Let $\Omega$ be an event space and $T$ a finite time horizon. We consider a $d$-dimensional ($d\geq 1$) stochastic process $\mathbf{x}(\omega,t):\Omega\times[0,T]\mapsto\mathbb{R}^{d}$ driven by an unknown (non-autonomous) stochastic differential equation (SDE) subject to external inputs
where $\mathbf{W}_{t}$ is $m$-dimensional ($m\geq 1$) Brownian motion, $\mathbf{a}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ drift function, $\mathbf{b}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d\times m}$ diffusion function, and $\mu(t)$ and $\nu(t)$ time-dependent external inputs into the stochastic system. In practice, the inputs can be external excitation signals or control signals. Throughout this paper, we will generally refer them as excitations and denote $\mathbf{u}(t)=(\mu(t),\nu(t))^{T}$ to consolidate the notation.
Our basic assumption is that the SDE is unknown, in the sense that the functions $\mathbf{a}$ and $\mathbf{b}$ are not known. Also, the driving Brownian motion $\mathbf{W}_{t}$ can not be observed. However, we have input-output (I/O) time history data between the excitations $\mathbf{u}=(\mu,\nu)^{T}$, i.e., the inputs, and system response $\mathbf{x}$, i.e., the output,
@@ 2.1 Problem Statement
(3) $\left(\mathbf{u}\left(t_{0}^{(i)}\right),\mathbf{u}\left(t_{1}^{(i)}\right),\cdots,\mathbf{u}\left(t_{L_{i}}^{(i)}\right)\right)\rightarrow\left(\mathbf{x}\left(t_{0}^{(i)}\right),\mathbf{x}\left(t_{1}^{(i)}\right),\cdots,\mathbf{x}\left(t_{L_{i}}^{(i)}\right)\right),$
@@ 2.2 Related Work and Contribution
This method developed in this paper has its foundation in two recent work: flow map learning (FML) for modeling deterministic unknown dynamical systems and its extension to modeling stochastic dynamical systems.
For an unknown deterministic autonomous system, $\frac{d\mathbf{x}}{dt}=\mathbf{f}(\mathbf{x})$, $\mathbf{x}\in\mathbb{R}^{d}$, where $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is unknown. The FML method seeks to approximate the unknown flow map $\mathbf{x}_{n}=\mathbf{\Phi}_{t_{n}-t_{s}}(\mathbf{x}_{s})$ by using observation data. More specifically, by using data on $\mathbf{x}$ over one time step $\Delta t$, the FML method constructs a model
@@
The primary contribution of this paper is on the development of data driven modeling for unknown stochastic systems subject to external excitations. To accomplish this, we extend the sFML framework ([10]), which was developed for autonomous system, to non-autonomous stochastic system. To learn the system I/O responses, we employ the local parameterization technique developed for non-autonomous deterministic system ([36]). The method parameterizes the input excitations in the data and transforms the learning problem into learning a parametric dynamical system. For stochastic non-autonomous system considered in this paper, we incorporate the method into a generative model in the sFML framework. In particular, we use normalizing flow as the generative model, which has not been considered in stochastic dynamical system learning. We shall demonstrate that the newly developed method is highly effective in modeling unknown stochastic systems, when excitations are not present in the training data.
@@ 3 Methodology
In this section, we describe the proposed learning method in detail.
@@ 3.1 Parameterization of Inputs
Consider the unknown SDE (1) over a time interval $[t_{n},t_{n+1}]$, $n\geq 0$,
@@
which can be wrriten equivalently as,
@@
are the expansion coefficients. In principle, one can choose any suitable basis functions. Since the time interval $[t_{n},t_{n+1}]$ usually has a (very) small step size $\Delta$, it suffices to use low-order polynomials. In fact, low-degree monomials bases, $p_{k}(\tau)=\tau^{k-1}$, $k\geq 1$, would be sufficient for most problems. When $k=0$, the parameterization takes form of piecewise constant function; when $k=1$, piecewise linear function.
The local parameterization of $\mathbf{u}$ is carried out based on the information one has about the excitations. If the excitations are only known at the discrete time instances, as shown in (3), then it is natural to utilize piecewise linear polynomial,
@@
If more information about $\mathbf{u}(t)$ is available, one can construct a higher degree polynomial. Note that since the time step $\Delta$ is usually small, a quadratic polynomial $\widetilde{\mathbf{u}}$ can be highly accurate in any time interval $[t_{n},t_{n+1})$. We remark that in the representation, only the values of the excitations $\mathbf{u}$ at $t_{n}$ and $t_{n+1}$ are needed. The values of the time $t_{n}$ and $t_{n+1}$ are not required.
@@ 3.2 Parametric Stochastic Flow Map
where the excitation signals $\mathbf{u}=(\mu,\nu)^{T}$ has been parameterized by $\widetilde{\mathbf{u}}$ via a set of parameters $\mathbf{\Gamma}_{n}$. Compared to (6), the transformed system (9) contains possible numerical error introduced by the parameterization of the excitations over the time domain $[0,\Delta)$. The error can be made arbitrarily small if one uses higher degree polynomials when $\Delta$ is sufficiently small.
By using subscript to denote the time level and letting
@@
the parameterized system (9) indicates that, there exists a mapping
@@ 3.3.1 Training Data
To construct the training data set, we reorganize the original training data set (3) into pairs of consecutive time instances. Since for each of the $i$-th trajectory, $i=1,\dots,N_{T}$, we can extract $L_{i}$ such pairs, there are a total number of $M=L_{1}+\cdots+L_{N_{T}}$ I/O data pairs from the data set (3):
@@ 3.3.2 Generative Model
and utilize the training data set (13) to learn the unknown operator $\mathbf{\widehat{G}}$. Note that the random variable $\mathbf{z}_{0}$ is not in the training date set. In practice, one chooses $\mathbf{z}_{0}$ with a known distribution, typically a standard Gaussian, and a specified dimension $n_{z}\geq 1$. The presence of the random variable $\mathbf{z}_{0}$ enables (15) to be a stochastic generative model that can produce random realizations. Several methods exist to construct stochastic generative models, e.g., GANs, diffusion model, normalizing flow, autoencoder-decoder, etc. In this paper, we adopt normalizing flow for (15).

### v2

@@ Abstract
We present a numerical method for learning an unknown nonautonomous stochastic dynamical system s, i.e., stochastic system s subject to time dependent excitation or control signals. Our basic assumption is that the governing equations for the stochastic system are unavailable. However, short bursts of input/output (I/O) data consisting of certain known excitation signals and their corresponding system responses are available. When a sufficient amount of such I/O data are available, our method is capable of learning the unknown dynamics and producing an accurate predictive model for the stochastic responses of the system subject to arbitrary excitation signals not in the training data. Our method has two key components: (1) a local approximation of the training I/O data to transfer the learning into a parameterized form; and (2) a generative model to approximate the underlying unknown stochastic flow map in distribution. After presenting the method in detail, we present a comprehensive set of numerical examples to demonstrate the performance of the proposed method, especially for long-term system predictions.
@@ 1 Introduction
There has been a growing interest in recovering/discovering unknown dynamical systems from observational data. Most of the existing studies focus on deterministic systems, with methods such as physics-informed neural networks (PINNs) [48, 49], SINDy [7], Fourier neural operator (FNO) [30], computational graph completion [39], sparsity promoting methods [50, 51, 24], flow map learning (FML) [47, 13], to name a few.
@@
Learning unknown stochastic systems is notably more challenging, as the stochastic noises in the systems usually can not be directly observed. The existing work utilizes Gaussian process [58, 2, 15, 37], polynomial approximations [54, 28], deep neural networks (DNNs) [10, 57, 11, 60, 17, 61], Koopman operators [14, 34], etc. More recently, a stochastic extension of the deterministic flow map learning (FML) approach [47, 13] was proposed. It employs generative models such as GANs (generative adversarial networks) [12] or autoencoders [56] to model the underlying stochasticity. However, most, if not all, of these methods are developed for autonomous systems, where time-invariance (in distribution) holds true and is critical to the method development.
@@
The focus and contribution of this paper is on the learning and modeling of unknown non-autonomous stochastic systems. More specifically, we consider SDEs with unknown governing equations and subject to time dependent external excitation or control signals. Our goal is to develop a method that can capture the stochastic dynamics of the unknown systems by using short-term data consisting of input/output (I/O) relations between the excitation signals and their corresponding system responses. We remark that there exist some studies on modeling deterministic non-autonomous systems, using methodology such as Dynamic Mode Decomposition (DMD) [31, 43, 26, 23]), SINDy ([8], Koopman operator [44, 33, 38], FML [45], etc. These methods are not applicable for stochastic non-autonomous systems.
@@
The proposed method in this paper has two key components. First, our method utilizes the observational I/O data to construct an accurate representation of unknown stochastic dynamics of the system. This is accomplished by a generative model that learns the stochastic mapping of the system between two consecutive discrete time steps. The learning of this stochastic flow map is similar to the work of [12, 56], which extended the deterministic FML to stochastic systems. While [12, 56] utilized GANs and autoencoder as the generative model, in this paper we employ conditional normalizing flow (cf. [40]). Normalizing flow has been widely adopted as a probabilistic model for generating data with desired distributions. Its applications include image and video generation [27], statistical inference and sampling [35, 52], reinforcement learning [22], as well as scientific computing [32, 29, 21, 16]. It has also been used to model temporal evolution of density function for stochastic system [59]. In this paper, we demonstrate that normalizing flow is applicable to modeling unknown stochastic systems and can serve as an effective alternative. The second key component of the proposed method is local parameterization of the excitation signal in the training I/O data. The local parameterization approach was first introduced in [45] for deterministic nonautonomous sytem. We adopt the similar idea and extend it to stochastic system. The approach seeks to parameterize the excitation sigals in the training data via a local polynomial over one time step, and subsequently transform the non-autonomous stochastic system into a local Markovian system with homogeneous increments, provided that the unknown stochastic input has stationary and independent increments. This in turn transforms the learning problem into a parametric learning between the coefficients of the local polynomials and the system responses. This is a critical component, as it allows the learned system to conduct long-term system predictions under arbitrary excitation signals that are never seen in the training data. Although the proposed method requires a large number of short bursts data, the overall demand for long trajectory data may not be as large, because each burst of the training I/O data consist of merely two time steps of data entries. Once trained, the learned model is able to simulate the unknown stochastic systems for very long time and subject to arbitrary exitation/control signals. A prominent feature of the method, perhaps unique to most of the existing methods, is that it does not require the unknown system to be a classical SDE (with drift and diffusion) and is applicable to learning general non-Gaussian stochastic systems.
@@ 2 Setup
Let $\Omega$ be an event space and $T$ a finite time horizon. We consider a $d$-dimensional ($d\geq 1$) stochastic process $\mathbf{x}(\omega,t):\Omega\times[0,T]\mapsto\mathbb{R}^{d}$ driven by an unknown non-autonomous stochastic dynamical system in the following general form
(1) $\frac{d\mathbf{x}_{t}}{dt}(\omega)=\mathbf{f}(\mathbf{x}_{t},\mathbf{u}(t),\omega),$
where $\omega\in\Omega$ represents the random input and $\mathbf{u}(t)$ are external excitation or control signals. Our basic assumption in this paper is that the governing equations of the stochastic system are not available. That is, the right-hand-side (RHS) $\mathbf{f}$ is unknown. Consequently, the solution $\mathbf{x}_{t}$ can not be obtained by solving (1).
We assume obervation data are available. More specifically, we have input-output (I/O) time history data between the excitations $\mathbf{u}$, i.e., the inputs, and system response $\mathbf{x}$, i.e., the output,
@@
Example 1: As an example of the unknown stochastic system (1), let us consider a non-autonomous stochastic differential equation (SDE)
where $\mathbf{W}_{t}$ is $m$-dimensional ($m\geq 1$) Brownian motion, $\mathbf{a}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d}$ drift function, $\mathbf{b}:\mathbb{R}^{d}\times\mathbb{R}\to\mathbb{R}^{d\times m}$ diffusion function, and $\mu(t)$ and $\nu(t)$ time-dependent external inputs into the stochastic system. In our setting, both the drift $\mathbf{a}$ and the diffusion $\mathbf{b}$ are unknown, and the excitation signal is $\mathbf{u}(t)=(\mu(t),\nu(t))^{T}$. Also, the driving Brownian motion $\mathbf{W}_{t}$ can not be observed. We emphasize that even though many stochastic system can be modeled in the form of (3), our proposed method does not require such a structure. In fact, the proposed method works for unknown stochastic system of the general form (1), where the random inputs do not need to be Gaussian or Brownian motion.
Example 2: Another example is reaction network, which can be modeled as stochastic jump process. For example, consider a gene expression model in the form of biochemical reaction network [53, 6],
(4) $\begin{split}\emptyset&\xrightarrow{k(t)}M,\\ M&\xrightarrow{k_{\text{s}}}M+P,\\ M&\xrightarrow{k_{\text{dm}}}\emptyset,\\ P&\xrightarrow{k_{\text{dp}}}\emptyset,\end{split}$
where $M$ is mRNA population and $P$ is protein population, and paramters $k$’s are reaction rates. Due to external excitation, one of the reaction rates becomes time dependent $k(t)$, whereas the other reaction rates remain constant. The dynamical behavior of $M$ and $P$ can be modeled by an exact Monte Carlo method known as Gillespie algorithm, or stochastic simulation algorithm (SSA), cf. [20]. In our setting, we assume that the reaction network (4) are not known. Subsequently, it is not possible to conduct SSA to predict the system behavior. We assume we can collect time history data of $(M,P)$ subject to certain known external signal $k(t)$. That is, in the form of the I/O data (2), the state variables are $\mathbf{x}(t)=(M,P)^{T}$ and the excitation signal is $\mathbf{u}(t)=k(t)$. Our goal is to model the long-term dynamics of $\mathbf{x}(t)$ as a function of $k(t)$. Note that the stochastic process $\mathbf{x}(t)$ can not be modeled by the SDE (3). It is not driven by Brownian motion and does not follow Gaussian distribution. This will be one of the numerical examples in the paper.
@@ 2.1 Problem Statement
(5) $\left(\mathbf{u}\left(t_{0}^{(i)}\right),\mathbf{u}\left(t_{1}^{(i)}\right),{\color[rgb]{0,0,0}\ldots},\mathbf{u}\left(t_{L_{i}}^{(i)}\right)\right)\rightarrow\left(\mathbf{x}\left(t_{0}^{(i)}\right),\mathbf{x}\left(t_{1}^{(i)}\right),{\color[rgb]{0,0,0}\ldots},\mathbf{x}\left(t_{L_{i}}^{(i)}\right)\right),$
@@ 2.2 Related Work and Contribution
This method developed in this paper has its foundation in two recent works: flow map learning (FML) for modeling deterministic unknown dynamical systems and its extension to modeling stochastic dynamical systems.
For an unknown deterministic autonomous system, $\frac{d\mathbf{x}}{dt}=\mathbf{f}(\mathbf{x})$, $\mathbf{x}\in\mathbb{R}^{d}$, where $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is unknown , the FML method seeks to approximate the unknown flow map $\mathbf{x}_{n}=\mathbf{\Phi}_{t_{n}-t_{s}}(\mathbf{x}_{s})$ by using observation data. More specifically, by using data on $\mathbf{x}$ over one time step $\Delta t$, the FML method constructs a model
@@
The primary contribution of this paper is on the development of data driven modeling for unknown stochastic systems subject to external excitations. To accomplish this, we extend the sFML framework ([12]), which was developed for autonomous system, to non-autonomous stochastic system. To learn the system I/O responses, we employ the local parameterization technique developed for non-autonomous deterministic system ([45]). The method parameterizes the input excitations in the data and transforms the learning problem into learning a parametric dynamical system. For stochastic non-autonomous system considered in this paper, we incorporate the method into a generative model in the sFML framework. In particular, we use normalizing flow as the generative model, which has not been considered in stochastic flow map learning. We shall demonstrate that the newly developed method is highly effective in modeling unknown stochastic systems, when excitations are not present in the training data.
@@ 3 Methodology
In this section, we describe the proposed learning method in detail. Since it is difficult to present the mathematical content in the general form (1), without knowledge of the equation and the stochastic inputs, our exposition in this section is primarily base on the more concrete SDE setting (3). It shall be clear that the proposed method does not rely on the special structure provided by (3).
@@ 3.1 Parameterization of Inputs
Again, for the convenience of mathematical exposition, let us consider the unknown SDE (3) over time interval $[t_{n},t_{n+1}]$, $n\geq 0$,
@@
which can be written equivalently as,
@@
are the expansion coefficients. In principle, one can choose any suitable basis functions. Since the time interval $[t_{n},t_{n+1}]$ usually has a (very) small step size $\Delta$, it suffices to use low-order polynomials. In fact, low-degree monomials bases, $p_{k}(\tau)=\tau^{k-1}$, $k\geq 1$, would be sufficient for most problems. When $k=0$, the parameterization takes form of piecewise constant function; when $k=1$, piecewise linear function. If higher degree polynomial approximation is needed, one should employ orthogonal polynomials as the basis for numerical stability.
The local parameterization of $\mathbf{u}$ is carried out based on the information one has about the excitations. If the excitations are only known at discrete time instances, as shown in (5), then it is natural to utilize piecewise linear function,
@@
If more information about $\mathbf{u}(t)$ is available, one can construct a higher degree polynomial. We remark that in the representation, only the values of the excitations $\mathbf{u}$ at $t_{n}$ and $t_{n+1}$ are needed. The values of the time $t_{n}$ and $t_{n+1}$ are not required.
It will become clear in the following sections that the critical approximation in the proposed method is the approximation in distribution (6), which is accomplished by a stochastic generative model via the observation data (5). Here, the numerical errors induced by randomness and finite sampling are dominant and overwhelm the small deterministic error of the local polynomial approximation. We have found no numerical advantage of using polynomials beyond second order.
@@ 3.2 Parametric Stochastic Flow Map
where the excitation signals $\mathbf{u}=(\mu,\nu)^{T}$ has been parameterized by $\widetilde{\mathbf{u}}$ via a set of parameters $\mathbf{\Gamma}_{n}$. Compared to (8), the transformed system (11) contains possible numerical error introduced by the parameterization of the excitations over the time domain $[0,\Delta)$. The error can be made arbitrarily small for smooth excitations if one uses higher degree polynomials when $\Delta$ is sufficiently small.
By using subscripts to denote the time level and letting
@@
the parameterized system (11) indicates that there exists a mapping
@@
Remark 3.2.
The derivation of (12), via the suppression of the time variable $t_{n}$, represents a subtle and yet signficant step. The parametric stochastic flow map (12) is Markovian, in the sense that $\widetilde{\mathbf{x}}_{n+1}$ depends only on the system information at $t_{n}$, i.e., on $\widetilde{\mathbf{x}}_{n}$, the local parameters for the excitation signal $\boldsymbol{\Gamma}_{n}$, and the stationary local stochastic increment $d\mathbf{W}_{n}$.
It is also clear that the derivation of (12) does not rely on the special structure of (3), so long as the stochastic input has stationary and independent increments. Therefore, we conclude that the parametric stochastic flow map (12) is applicable to the general system (1) when the stochastic inputs are Lévy processes.
@@ 3.3.1 Training Data
To construct the training data set, we reorganize the original training data set (5) into pairs of consecutive time instances, often referred to as snapshots. Since for each of the $i$-th trajectory, $i=1,\dots,N_{T}$, we can extract $L_{i}$ such pairs, there are a total number of $M=L_{1}+\cdots+L_{N_{T}}$ I/O data pairs from the data set (5):
@@ 3.3.2 Generative Model
and utilize the training data set (15) to learn the unknown operator $\mathbf{\widehat{G}}$. Note that the random variable $\mathbf{z}_{0}$ is not in the training date set. In practice, one chooses $\mathbf{z}_{0}$ with a known distribution, typically a standard Gaussian, and a specified dimension $n_{z}\geq 1$. The presence of the random variable $\mathbf{z}_{0}$ enables (17) to be a stochastic generative model that can produce random realizations. Several methods exist to construct stochastic generative models, e.g., GANs, diffusion model, normalizing flow, autoencoder-decoder, etc. In this paper, we adopt normalizing flow for (17). We remark that the use of normalizing flow is not critical to the applicability of the proposed method. Our numerical experiments have indicated that other generative models such as GANs and diffusion model work generally well. We employ normalizing flow primarily because it has not been used in the FML framework before. The property, applicability and performance comparison of different generative models is a rather complex issue and out of the scope of this paper.